OpenStudy (hellokitty17):

The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

oregonduck (oregonduck):

@HootHoot

OpenStudy (hoothoot):

lol

OpenStudy (hoothoot):

hen your stuck on a question who u cal hoothoot

OpenStudy (hellokitty17):

lol

oregonduck (oregonduck):

or oregonduck because a) Let s represent the side of the square. Then you have .. x + 1 = s = y - 5 ... where y is the length .. y = x+6 ... solve for y Now the area is the product of width (x) and length (y) .. x(x+6) = 280 b) x^2 + 6x + 9 = 289 ... add 9 to complete the square .. (x+3)^2 = 17^2 ... rewrite as squares .. x = 14 ... take the square root and subtract 3 The width of the land is 14 meters.

OpenStudy (hoothoot):

hen the length is 5 less, and the width is 3 more, then both dimensions are equal `because it's a square`.\[\Large\bf\sf \ell-5\quad=\quad w+3\]

OpenStudy (hoothoot):

ok hellokity relie on mr duck my side kick lol to give answer im tired im a go to bed